OpenStudy (anonymous):

The 3 sides of a right triangle are x, x+5, and 2x-5 (hypotenuse). Find x. Please help!

OpenStudy (anonymous):

@Agl202

OpenStudy (anonymous):

have you ever heard of the term a^2+b^2=c^2?

OpenStudy (anonymous):

yes, pythagorean therom

OpenStudy (anonymous):

x^2 + (x+5)^2=(2x-5)^2

OpenStudy (anonymous):

Have you tried plugging it into that and solving for x? And yes it will take a lot of work

OpenStudy (anonymous):

now expand all the squares

OpenStudy (anonymous):

What do you mean by expanding the squares?

OpenStudy (anonymous):

ok so when you have the equation (x+5)^2 its the same as (x+5)(x+5) and when you multiply it out using FOIL you get an equation and you combine like terms

OpenStudy (anonymous):

so for x would it still be x^2 after foiling it? x+5 would be x^2 +10x+25

OpenStudy (anonymous):

Correct! no try the other side :)

OpenStudy (anonymous):

now not no

OpenStudy (anonymous):

k so that would be um... 4x-20x+25

OpenStudy (anonymous):

Is that correct?

OpenStudy (anonymous):

Correct! now combine your like terms on the left side then subtract your left side from both sides to get one equation and to get your x on one side

OpenStudy (anonymous):

x^4-6x?

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

Now I just need to find x

OpenStudy (anonymous):

yep :) now can you factor anything out of that?

OpenStudy (anonymous):

just x i think

OpenStudy (anonymous):

hold on hun how did you get a 4th power on the x?

OpenStudy (anonymous):

x^2 times x^2

OpenStudy (anonymous):

when you combined on the left side all you do is add coefficients whenever you multiply then you add the powers together

OpenStudy (anonymous):

nevermind i made a mistake lol no ones perfect :P

OpenStudy (anonymous):

so what would it be instead of x^4?

OpenStudy (anonymous):

but yes your right you can only factor an x out so that leaves you with x(x^3-6)

OpenStudy (anonymous):

ok so how do i figure out what x is from that? I need to put in an exact length in feet, it cannot be an equation. @goatsgirl17

OpenStudy (anonymous):

well = the on in the equation to zero then add 6 to both sides and find the cubed route of that and the other x equals 0

OpenStudy (anonymous):

im confused, how do i do that?

OpenStudy (anonymous):

@goatsgirl17

OpenStudy (anonymous):

x^3=6 therefore type into the calculator 2nd ^ then type in 3 before the bar and put a 6 under the bar

OpenStudy (anonymous):

bar?

OpenStudy (anonymous):

the square root im sorry i gtg

OpenStudy (anonymous):

ok thanks